### v3

@@ Abstract
Airdrops are a common strategy used by blockchain protocols to attract and grow an initial user base. Tokens are typically distributed to select users as a “reward” for engaging with the protocol, aiming to foster long-term community loyalty and sustained economic activity. Despite their prevalence, there is limited understanding of what makes an airdrop successful. This paper outlines the design space for airdrops and proposes key outcomes for an effective strategy. We analyze on-chain data from six large-scale airdrops to assess their success and find that a substantial portion of tokens is often sold off by “airdrop farmers.” Based on this analysis, we highlight common pitfalls and offer guidelines for improving airdrop design.
1 Introduction
Blockchain protocols often design reward programs to attract new customers and foster the loyalty of existing users. In recent years, the practice of distributing platform-minted tokens, known as “airdropping,” has gained widespread popularity [45]. For example, in 2023 alone, 4.56 billion US dollars worth of tokens were airdropped to users across various protocols [6]. Although the use of airdrops is prevalent in the blockchain world, we present preliminary findings showing that there is no significant correlation between performing an airdrop and a platform’s popularity relative to available alternatives. Intuitively, this is not optimal and may result in the loss of funds that can be used to improve the Quality-of-Service (QoS)offered by the platform.
@@ 1 Introduction
To understand why previous airdrops did not always achieve their goals and to quantify their success, we first articulate a set of reasonable desired outcomes for airdrops. We go over previous airdrops, see how they measure up, and uncover some interesting insights in the process, comparing them to a basic desiderata. Specifically, we analyze the data pertaining to five popular airdrops (ENS [14], dYdX [33], 1inch [7], Arbitrum [35], and Uniswap [1]) and a fake airdrop conducted by a Sybil farmer (Gemstone [30, 42]). We conclude that most of the proceeds (as high as 95%) have been sold via exchanges shortly after the airdrop, implying that the airdrops fall short of their intended purpose and benefit mostly airdrop farmers, highly-skilled users who employ sophisticated tactics to increase the share of tokens that they receive. Furthermore, we describe the common challenges faced by airdrops in the past. Given that airdrops are a relatively recent phenomenon, it is expected that the theoretical and practical understanding of them are still in their infancy, and thus previous airdrops may not have been entirely successful in the long run. Finally, we use our analysis to suggest ways to create better airdrop mechanisms that are more fair to honest users.
@@
Quantitative analysis. We conducted a quantitative analysis of airdrops executed by ENS, dYdX, 1inch, Arbitrum, Uniswap, and an additional fake airdrop named Gemstone. We show that most of the funds received through these airdrops are sold on exchanges, instead of being used to create dapp or platform engagement on the part of end-users. This corresponds to 36.62% of all ENS tokens. Similar trends are evident for dYdX (35.45%) and 1inch (54.05%). On average, tokens were typically sold on exchanges within 1 to 2.34 transfers after being received, with a median of 2 transfers.
Qualitative analysis. We conducted a qualitative analysis of past airdrops and propose guidelines for the development of future airdrops to mitigate the issues we identified. Our focus lies on airdrop farming and the distribution of governance tokens through airdrops. To address these concerns, we suggest alternative incentives, such as offering fee discounts for subsequent interactions within the blockchain protocol.
Multi-chain empirical data. We collected data from two major Ethereum rollups that are underexplored by the literature, Arbitrum and ZKsync Era, and moreover, collect and label data from major airdrops. We plan to share both our dataset and scripts in a publicly-accessible repository.
@@ 2 Airdrop Desiderata
Moreover, the timing of an airdrop significantly impacts the number of eligible users receiving tokens. Established protocols opting for delayed airdrops may have a larger user base compared to newer projects executing airdrops sooner. This difference in user base size introduces complexity when performing Sybil detection, as there are more accounts to evaluate and potentially safeguard against. If not performed correctly, this could impact community happiness. In such instances, the protocol may inadvertently reward accounts linked to industrial farming, which could be viewed negatively by the community. To mitigate this issue, LayerZero Labs implemented a self-reporting mechamism for sybils [39]. Under this system, sybils have the option to self-report and receive 15% of their designated token allocation.
Airdrop | Start date | End date | Blockchain | Occurrence | Project type |
ENS [14] | Oct. 2021 | May 2022 | Ethereum | Single | Name service |
dYdX [33] | Sept. 2021 | Still active | Ethereum | Continuous | DEX |
1inch [7] | Dec. 2020 | Still active | Ethereum | Continuous | DEX |
Gemstone [30] | Jul. 23, 2023 | Jul. 31, 2023 | ZKsync Era | Single | Fake airdrop |
Arbitrum [35] | Mar. 2023 | Sept. 2023 | Arbitrum One | Single | Single |
Uniswap [1] | Sept. 2020 | Still active | Ethereum | Continuous | DEX |
@@
Target value-creating users. An airdrop should ideally focus on users who can contribute the most to the long-term sustainability of their platform. In the context of platforms that rely on user-based liquidity provisioning, this may refer to users who provide the most liquidity to lending pools [72] and Decentralized Exchanges (DEXs) [70], or across a variety of tokens. In rollups, one may consider so-called “creators” who deploy popular and useful contracts or users who bridge tokens to the rollup [74] as particularly valuable. Such users provide additional use cases for a platform, and thus serve to attract other users.
@@ 3 Exchange Aftermarket Analysis
The motivation for our quantitative analysis in this paper comes from the following observation: airdrop recipients quickly sell tokens and jump ship, which obviously defeats the point of airdrops in the first place. Analyses performed on DEXairdrops show that airdrop recipients sometimes sell all of their tokens shortly after receiving them. For example, shortly after the ParaSwap’s drop, 61% of their tokens where sold [16].
In both cases, the majority of recipients ceased using the relevant blockchain protocols within a few months. This pattern suggests that the airdrop was ineffective in keeping recipients engaged, or that many recipients were likely Sybil accounts. Moreover, a rapid sell-off could unsettle the market, particularly if seen as an indication of dwindling confidence in the protocol’s future prospects. Here, we examine data associated with six airdrops presented in Table 1, which were gathered from archive nodes of Ethereum, Arbitrum, and ZKsync Era (see §0.A for details on our data collection). To identify exchanges, we use a list of 620 exchange account addresses obtained from Dune and Etherscan [15].
@@
ENS | 36.62% | |||||||
dYdX | 35.45% | |||||||
1inch | 54.05% | |||||||
Gemstone | 95.01% | 0 | ||||||
Arbitrum | 10.15% | |||||||
Uniswap | 5.50% |
3.1 Tokens Distributed Over Time
We provide quantitative insights into how airdropped tokens are disseminated across the user base in Table 2. Our analysis reveals a significant number of recipients for each airdrop, suggesting potential involvement of airdrop farmers (see column 5 in Table 2). Furthermore, the data indicate that tokens were frequently traded on exchanges. This is supported by the observation that the initial transfer made by the addresses post-airdrop was often a sale on an exchange (see column 4 in Table 2). Notably, the case of Gemstone is particularly striking, with 95% of its tokens being sold on exchanges. In this particular Gemstone case, a non-open sourced DEX created by the airdrop farmer [30].
Furthermore, Gemstone’s airdrop stands out significantly in terms of the number of tokens distributed compared to other airdrops. This substantial distribution also results in a notably higher median number of airdrops per recipient when compared to the other airdrops as shown in Table 2. Remarkably, Gemstone allocated the lion’s share, specifically 99.53%, of their total token supply during their airdrop campaign. It is important to highlight that Gemstone was executed primarily as a Sybil attack rather than a legitimate airdrop.
Among all the airdrops analyzed in this study, the Arbitrum airdrop exhibited the fastest rate of token claims by users. Fig. 1(a) illustrates the distribution of accounts claiming their tokens each day. Arbitrum conducted a substantial airdrop, distributing ARB tokens among selected accounts. Notably, (93.28%) of these accounts successfully claimed 94.03% of the allocated ARB tokens. Remarkably, 72.45% of the accounts completed their token claims on the first day, with an additional 14.41% doing so on the following day. This cumulative figure amounts to almost 87% of accounts claiming the Arbitrum airdrop within one day of its release, indicating high user engagement and prompt action among the majority of participants.

### v4

@@ Abstract
Airdrops are a popular mechanism used by blockchain protocols to bootstrap communities, reward early adopters, and decentralize token distribution. Despite their widespread adoption, the effectiveness of airdrops in achieving long-term user engagement and ecosystem growth remains poorly understood. In this paper, we present the first comprehensive empirical study of nine major airdrops across Ethereum and Layer-2 ecosystems. Our analysis reveals that a substantial share of tokens—up to 66% in some cases—are rapidly sold, often in recipients’ first post-claim transaction. We show that this behavior is largely driven by “airdrop farmers,” who strategically optimize eligibility criteria to extract value without contributing meaningfully to the ecosystem. We complement our quantitative findings with a case study of the Arbitrum airdrop, illustrating how short-term activity spikes fail to translate into sustained user involvement. Based on these results, we discuss common design pitfalls—such as Sybil vulnerability, poor incentive alignment, and governance token misuse—and propose actionable guidelines for designing more effective airdrop strategies.
keywords:
Airdrop design, Airdrop farmer, Web3, Layer 21 Introduction
Blockchain protocols often design reward programs to attract new customers and foster the loyalty of existing users. In recent years, the practice of distributing platform-minted tokens, known as “airdropping,” has gained popularity [56]. For example, in 2023 alone, 4.56 billion US dollars worth of tokens were airdropped by various protocols [9].
Although the use of airdrops is prevalent in the blockchain world, we present preliminary findings showing that there is no significant correlation between performing an airdrop and a platform’s popularity relative to available alternatives. Intuitively, this is not optimal and may result in the loss of funds that can be used to improve the Quality-of-Service (QoS)offered by the platform.
@@
To understand why previous airdrops did not always achieve their goals and to quantify their success, we first articulate a set of reasonable desired outcomes for airdrops. We go over previous airdrops, see how they measure up, and uncover some interesting insights in the process, comparing them to a basic desiderata. To ground our analysis in real-world observations, we focus on nine widely recognized airdrops: 1inch [10], Arbitrum [45], Arkham [5], DYDX [43], ENS [20], Lido [51], Optimism [31], Tornado Cash [69], and Uniswap [1]. These protocols were selected to ensure diversity across blockchain ecosystems (including Ethereum and Layer-2 rollups), protocol categories (e.g., DEXs, governance platforms, privacy tools, staking services), and market capitalizations (refer to Table 1). This curated selection allows us to draw generalizable insights while maintaining analytical depth across a representative cross-section of the blockchain landscape.
@@
We conclude that most of the proceeds (as high as 65.75%) have been sold via exchanges shortly after the airdrop, implying that the airdrops fall short of their intended purpose and benefit mostly airdrop farmers, highly skilled users who employ sophisticated tactics to increase the share of tokens that they receive. One notable example of such farming activities occurred during the ZKsync Era. In this incident, a Sybil farmer orchestrated a fake airdrop, utilizing over accounts and deploying a fake, non-open-sourced DEX to qualify for the ZKsync token airdrop [38, 53]. We analyze this incident and discuss it in §A. Furthermore, we describe the common challenges faced by airdrops in the past. Given that airdrops are a relatively recent phenomenon, it is expected that the theoretical and practical understanding of them are still in their infancy, and thus previous airdrops may not have been entirely successful in the long run. Finally, we use our analysis to suggest ways to create better airdrop mechanisms that are fairer to honest users.
1.1 Research Questions
We address three critical research questions central to our study.
RQ-1: Do airdrops lead to long-term user retention?
Airdrops are widely used to bootstrap user bases by distributing tokens to early adopters. However, their long-term effectiveness remains questionable. We examine whether protocols that conducted airdrops successfully retained users beyond the initial distribution period.
Our analysis shows that while airdrops temporarily increase user activity, they often fail to ensure lasting engagement. For instance, even widely publicized airdrops such as those conducted by Arbitrum or Optimism experienced an initial surge in user participation followed by a noticeable decline. These patterns are consistent with external reports, such as those by Messari [8], which observe that users often migrate to protocols offering new or better incentives once the airdrop rewards are exhausted.
Interestingly, user migration is not limited to protocols that conduct airdrops. Platforms without current airdrops also attract users motivated by the expectation of future distributions. Our findings suggest that airdrops may be necessary to attract users initially, but insufficient on their own to build a committed, long-term community.
RQ-2: What alternative incentives can enhance airdrops to sustain user engagement?
Given the limitations of airdrops alone, we explore alternative mechanisms that could complement or extend their effectiveness. One promising strategy is to incentivize developers and builders to create applications such as Decentralized Exchanges (DEXs), Non-Fungible Token (NFT)platforms, and other dApps that enrich the ecosystem and enhance platform stickiness. This aligns user incentives with broader platform utility.
Protocols could also offer fee discounts or service credits to users holding airdropped tokens, creating ongoing utility for the asset beyond speculative value. Another viable approach is implementing multi-phase airdrops, allowing protocols to refine their criteria and distribution logic across rounds [2]. Finally, user-centric design such as incorporating community feedback into airdrop mechanisms can help protocols better align with user expectations and values.
However, these strategies must contend with user behavior patterns shaped by opportunism. Blockchain users are highly mobile and prone to chasing short-term gains. Without a compelling long-term value proposition, users are likely to disengage once rewards diminish.
RQ-3: What are the key design challenges faced when implementing airdrops?
Airdrop campaigns face several technical and economic challenges that can undermine their objectives. First, Sybil attacks remain a pervasive threat. Sophisticated airdrop farmers create pseudonymous wallets to game eligibility criteria and maximize token claims [23, 40, 41, 54, 62, 68, 80]. These actors may employ automated tools [11, 15, 32, 77] to mimic legitimate user behavior and evade Sybil-detection heuristics. Existing detection methods leverage graph analysis and machine learning [87, 54, 50], but the adversarial nature of these users makes defense an ongoing arms race.
Second, governance token distributions pose their own risks. When poorly designed, they may concentrate power in the hands of opportunistic or uninformed actors. Given that many governance systems rely on a one-token-one-vote mechanism, protocols risk empowering misaligned stakeholders [24, 59, 74]. This undermines governance integrity and can discourage community participation.
Finally, other issues, such as insider trading before an airdrop, preferential access to airdrop distribution criteria or leaks related to forthcoming airdrops can erode trust, as has been observed in several cases [25, 64, 75]. Ensuring transparency and fairness in airdrop design is therefore critical, not just for regulatory compliance, but for maintaining legitimacy in the eyes of the community.
1.2 Summary of our Contributions
@@
Quantitative analysis. We conducted a quantitative analysis of airdrops executed by 1inch, Arbitrum, Arkham, DYDX, ENS, Lido, Optimism, Tornado Cash, and Uniswap. We show that most of the funds received through these airdrops are sold on exchanges in the first airdropped token transfer made by the user after receiving the airdrop, instead of being used to create dapp or platform engagement on the part of end-users. This corresponds to 65.75% of all Lido tokens, followed by 1inch with 58.67%, and Optimism with 48.21%, among others. On average, tokens were typically sold on exchanges within 1.17 to 2.76 transfers after being received, with a median of 1 to 2 transfers.
Qualitative analysis. We conducted a qualitative analysis of past airdrops and propose guidelines for the development of future airdrops to mitigate the issues we identified. Our focus lies on airdrop farming, the distribution of governance tokens through airdrops, and insider trading. To address these concerns, we suggest alternative incentives, such as offering fee discounts for subsequent interactions within the blockchain protocol.
Multi-chain empirical data. We collected data from two major yet underexplored Ethereum rollups—Arbitrum and Optimism—as well as from nine major airdrops. We also gathered ZKsync Era data to discuss a fake airdrop performed by sybil farmers on ZKsync Era (see §A). Additionally, we curated address labels to identify centralized and decentralized exchange addresses. To promote reproducibility and support future research, we publicly release our dataset and data collection scripts in an open-access repository [60].
| Airdrop | First Claim Date | Last Claim Date | Blockchain | Occurrence | Project Type | Market Cap (USD) |
| 1inch [10] | Dec. 2020 | Still active | Ethereum | Continuous | DEX | $270.24M |
| Arbitrum [45] | Mar. 2023 | Sept. 2023 | Arbitrum One | Single | Rollup | $1.56B |
| Arkham [5] | Jul. 2023 | Still active | Ethereum | Continuous | Analytics | $118.89M |
| DYDX [43] | Sept. 2021 | Still active | Ethereum | Continuous | DEX | $68.71M |
| ENS [20] | Nov. 2021 | Aug. 2024 | Ethereum | Single | Name service | $586.96M |
| Lido [51] | Jan. 2021 | Still active | Ethereum | Continuous | Staking | $819.18M |
| Optimism [31] | May 2022 | Jul. 2023 | OP Mainnet* | Multi-round | Rollup | $1.24B |
| Tornado Cash [69] | Dec. 2020 | Aug. 2022 | Ethereum | Single | Mixer | $39.68M |
| Uniswap [1] | Sept. 2020 | Still active | Ethereum | Continuous | DEX | $3.9B |
@@ 2 Airdrop Desiderata
Moreover, the timing of an airdrop can impact the number of eligible users receiving tokens. Established protocols opting for delayed airdrops may have a larger user base compared to newer projects executing airdrops sooner. This difference in user base size introduces complexity when performing Sybil detection, as there are more accounts to evaluate and potentially safeguard against. If not performed correctly, this could impact community happiness. In such instances, the protocol may inadvertently reward accounts linked to industrial farming, which could be viewed negatively by the community. To mitigate this issue, LayerZero Labs implemented a self-reporting mechanism for sybils [49]. Under this system, sybils have the option to self-report and receive 15% of their designated token allocation.
@@
Target value-creating users. An airdrop should ideally focus on users who can contribute the most to the long-term sustainability of their platform. In the context of platforms that rely on user-based liquidity provisioning, this may refer to users who provide the most liquidity to lending pools [86] and DEXs [84], or across a variety of tokens. In rollups, one may consider so-called “creators” who deploy popular and useful contracts or users who bridge tokens to the rollup [88] as particularly valuable. Such users provide additional use cases for a platform, and thus serve to attract other users.
3 Data Collection
| Protocol | Address | Chain | Event type | Start block | End block | Start date | End date | Number of | ||
| Transfers | Claims | |||||||||
| 1inch | 0x1111c302 | Ethereum | Transfer | 2020-12-23 | 2024-10-19 | |||||
| Arbitrum | 0x912c6548 & 0x67a2c9d9 | Arbitrum | Transfer & Claim | 2023-03-16 | 2024-05-14 | |||||
| Arkham | 0x6e2ab050 & 0x08c7ab05 | Ethereum | Transfer & Claim | 2023-07-05 | 2024-10-19 | |||||
| DYDX | 0x92d6eff5 & 0x01d3a588 | Ethereum | Transfer & RewardsClaimed | 2021-07-12 | 2024-10-19 | |||||
| ENS | 0xc1839d72 | Ethereum | Transfer & Claim | 2021-11-01 | 2024-10-19 | |||||
| Lido | 0x5a981b32 | Ethereum | Transfer & ClaimedTokens | 2020-12-17 | 2024-10-19 | |||||
| Optimism | 0x42000042 & 0xfedf57de | Optimism | Transfer & Claim | 2022-04-26 | 2024-03-24 | |||||
| Tornado Cash | 0x7777116c | Ethereum | Transfer | 2020-12-18 | 2024-10-19 | |||||
| Uniswap | 0x1f98f984 | Ethereum | Transfer | 2020-09-14 | 2024-10-19 | |||||
To analyze airdrop distribution patterns and their post-distribution market behavior (§4), we compiled a comprehensive dataset from multiple blockchains and external data sources. Specifically, we collected data from Ethereum, Arbitrum, and Optimism using blockchain nodes, along with metadata from public sources such as Growthepie and address-labeling services.
On-chain data collection. Our primary dataset consists of smart contract event logs data, since their inception to October 2024, related to airdrop claims and token transfers for nine major protocols: 1inch, Arbitrum, Arkham, DYDX, ENS, Lido, Optimism, Tornado Cash, and Uniswap. For each project, we identified and extracted all relevant contract events, including airdrop eligibility, claim transactions, and subsequent token transfers. Table 2 lists the smart contracts and event types analyzed in this study. Some protocols—such as Arbitrum, Arkham, DYDX, and Optimism—use separate contracts for their usual token transfers and airdrop distribution. Others, including 1inch, ENS, Lido, Tornado Cash, and Uniswap, rely on a single contract to handle both token transfers and airdrop claims, typically distinguished by different event types (e.g., a Claim and Transfer events). In certain cases, such as 1inch, Tornado Cash, and Uniswap, the contract does not emit distinct events for Transfer and Claim. In these situations, we identify airdropped tokens by detecting transfers originating from known airdrop distribution addresses belonging to the respective projects.
Exchange and address labeling. To differentiate between regular users and exchange-associated addresses (both centralized and decentralized exchanges), we aggregated public address labels from Etherscan [22], Optimism Explorer [66], Arbiscan [4], Dune [39], and similar services. Additionally, we queried the Arkham Intelligence API [5] for high in-degree nodes—i.e., addresses that frequently receive tokens from many accounts—which are characteristic of exchange wallets. With these approaches, we labeled the top 150 such addresses for each of the nine protocols and ultimately identified a total of 3569 exchange addresses: 3011 centralized (CEX) and 558 decentralized (DEX).
Token distribution visualization. For the token distribution graphs presented in §4.2, we used the open-source tool Gephi to visualize address interactions and airdrop flows. To enhance the readability of the network visualizations, we focused on a subgraph consisting of each account’s first token transfer following the airdrop receipt.
Cross-chain ecosystem data. In §5, we complement our analysis with macro-level blockchain metrics from Growthepie.xyz [37] for the year 2023. These include daily transaction counts, active addresses, median transaction fees, Total Value Locked (TVL), user fees, and stablecoin market capitalizations across leading Layer-2 ecosystems (ZKsync Era, OP Mainnet, Arbitrum, Immutable X, and Polygon zkEVM). All time series were smoothed using a 7-day moving average to enhance interpretability.
Reproducibility. To ensure the reproducibility of our results, we publicly release the complete dataset and analysis scripts in an open-access repository [60].
@@ 4 Exchange Aftermarket Analysis
The motivation for our quantitative analysis in this paper comes from the following observation: airdrop recipients quickly sell tokens and jump ship, which obviously defeats the point of airdrops in the first place. Analyses performed on DEXairdrops show that airdrop recipients sometimes sell all of their tokens shortly after receiving them. For example, shortly after ParaSwap’s airdrop, 61% of their tokens were sold [23].
In both cases, the majority of recipients ceased using the relevant blockchain protocols within a few months. This pattern suggests that the airdrop was ineffective in keeping recipients engaged, or that many recipients were likely Sybil accounts. Moreover, a rapid sell-off could unsettle the market, particularly if seen as an indication of dwindling confidence in the protocol’s future prospects. In our study, we examine data associated with nine airdrops presented in Table 1, which were gathered from blockchain data of Ethereum, Arbitrum, Optimism, and ZKsync Era (as discussed in §3).
Interestingly, some addresses associated with DYDX, Arkham, Lido, and Tornado Cash received airdropped tokens multiple times. This could be attributed to fees, yield payments, or other protocol distributions. However, per Table 1, this is in line with their airdrop design that is continuous, meaning that the same address could qualify for the next airdrop round. In fact, DYDX addresses received an airdrop on average 1.58 times, with a standard deviation of 1.68, ranging from 1 to 30 instances, and a median of 1. For Arkham, five addresses received the airdrop twice. Similarly, nine addresses linked to Lido also received two airdrops. In the case of Tornado Cash, 99 addresses received airdrops more than once, with an average of 1.02 instances, a standard deviation of 0.26, a median of 1, and a range from 1 to 16 occurrences.
4.1 Tokens Distributed and Liquidated Over Time
We analyze how airdropped tokens were distributed and what actions recipients took following their receipt. We organize our findings into four key areas: claim timing, early liquidation behavior, interaction with exchanges, and continued protocol engagement.
4.1.1 Airdrop Claim Behavior
Figure 1 illustrates the rate at which eligible users claimed their airdropped tokens. The Arkham, Lido, and Arbitrum airdrops saw the fastest and most concentrated activity: 87%, 77%, 73%, respectively, of eligible accounts claimed tokens on the first day, and nearly 94%, 92%, and 87%, respectively, by the second day. This pattern suggests strong anticipation or automation, likely driven by airdrop farming tools or pre-announced eligibility windows. Other protocols, such as ENS, DYDX, and Uniswap, exhibited more gradual claiming behavior, with long tails spanning multiple weeks.
4.1.2 Initial Liquidation: Airdrops Sold Shortly After Receipt
Table 3 shows the percentage of tokens sold at exchanges as recipients’ first action post-claim (column 4). Lido had the highest liquidation rate (65.75%), followed by 1inch (58.67%) and Optimism (48.21%). In contrast, DYDX and Arkham experienced lower immediate liquidation rates at 13.56% and 19.79%, respectively. This variation likely stems from protocol-specific factors such as token utility, incentive structures, or community expectations.
These findings are supported by Figures 2 and 3, which reveal the timing and structure of token transfers. Specifically, Figure 3(a) shows that 66.09% of 1inch recipients transferred tokens to exchanges within 24 hours. Similarly, ENS and Lido saw over half of their recipients interact with exchanges in the same timeframe. Figure 3(b) indicates the results based on the number of blocks. The frequency at which blocks are committed varies according to the evaluated blockchain. That is, Ethereum processes blocks at a much lower pace than Arbitrum, Optimism, and ZKsync Era that are L2solutions.
@@ 4.1.2 Initial Liquidation: Airdrops Sold Shortly After Receipt
| 1inch | 58.67% | |||||||
| Arbitrum | 28.64% | |||||||
| Arkham | 19.79% | |||||||
| DYDX | 13.56% | |||||||
| ENS | 46.74% | |||||||
| Lido | 65.75% | 0.000871 | ||||||
| Optimism | 48.21% | |||||||
| Tornado Cash | 33.71% | |||||||
| Uniswap | 43.69% | |||||||
4.1.3 Exchange Interaction and Transfer Hops
To analyze the path tokens take from airdrop recipients to CEXs and DEXs, we consider a blockchain transaction graph and formally define the number of “hops” between a source node (i.e., an airdrop recipient) and a terminal node (i.e., an exchange).
Definition 4.1 (Number of hops to reach an exchange).
Let denote a graph where is the set of all blockchain addresses and an edge represents a transfer from address to address . Furthermore, let be the set of airdrop recipient addresses and be the set of exchange addresses. For any two addresses , denote by the length of the shortest path from to in . If no path exists, then .
Given an airdrop recipient , the number of hops to reach an exchange is defined as the shortest path from to any exchange in :
Table 4 quantifies how directly recipients sent tokens to exchanges. We find that most addresses liquidated tokens within 1–2 hops: for example, the median number of hops for Arbitrum, Arkham, and DYDX was 2, while others had medians of 1. The fraction of recipients who interacted with exchanges was substantial: over 60% for 1inch and Optimism, and over 75% for Lido. Only a few protocols, like Arkham (15.63%), had lower exchange interaction rates, likely due to token holding incentives or smaller allocations.
These interactions were not limited to simple sell-offs. As shown in Table 5, the median number of days to transfer to an exchange was almost under 1 for nearly all protocols. For instance, Arkham and DYDX saw medians of 0.01 and 0.04 days, respectively. Despite some outliers with longer delay times, the heavy left tail in Figure 3 indicates that a large majority of sales occurred almost immediately after claim.